OpenStudy (anonymous):

Find the slope and y-intercept y=3x+2

OpenStudy (anonymous):

slope = M plug 0 in for x to get y intercept y=mx+b

OpenStudy (anonymous):

I still don't get that rule...

OpenStudy (anonymous):

Slope Intercept form is y=mx+b so plug in 0. y=3(0)+2 and slop is M. Always

OpenStudy (anonymous):

So put 0 for x?

OpenStudy (anonymous):

Plug in 0 for x to get the y intercept